OpenStudy (anonymous):

If (a+b)^2 = 25, and "a" and "b" are negative integers, what is the value of (a+b)^3?

OpenStudy (anonymous):

Squaring makes the first one positive. So if you are multiplying 3 negatives, you get....? And since the answer to the first is 25, you know what a and b add up to, right?

OpenStudy (anonymous):

multiplying 3 negatives = positive because - times - = + and + times - = - and - times - = + so then (-5+-5)^3 = -1000 wait that's not right... the answer is supposed to be -125

OpenStudy (anonymous):

Ohhhhhh it's 2 and 3 not 5

OpenStudy (anonymous):

multiplying 3 negatives gets you a negative. I didn't say a and b were each -5. I said you know what they add up to be squared to 25. It doesn't matter what a and b are individually in this problem. So, since you know that a and b add to -5, then -5^3 is your answer.

OpenStudy (anonymous):

ok, i get it. was just misunderstanding what was being asked. is there another way of solving this?

OpenStudy (anonymous):

multiplying 3 negatives = positive because - times - = + and + times - = - and - times - = + And what you did here was multiply 4 negatives. :)

OpenStudy (anonymous):

Take the square root of both sides. (a+b)= +or - 5 (in this case we know it's -5 because they tell you a and b are both negatives. Then put it into the (a+b)^3=? equation.

OpenStudy (anonymous):

i don't understand what u mean by the first part. do u mean the square root of -5 or the 25? or

OpenStudy (anonymous):

The square root of (a+b) squared is just (a+b), right? And when you take the square root of something, the answer can either be positive or negative, so the square root of 25 is either +5 or -5.

OpenStudy (anonymous):

oh ohhhh, but how do u know which one to use, +5 or -5?

OpenStudy (anonymous):

Well, in this case, they told you that a and b are both negative numbers, so you know it HAS to be the -5. But in other cases, they could both be right. You read the details carefully. But to check, you just plug your answers back into the original equation and see if the equation is true. :)

OpenStudy (anonymous):

I suspect what they were really trying to get you to see from this problem is that if you cube a negative, you get a negative.

OpenStudy (anonymous):

i feel silly, ure right, i need to b more meticulous.

OpenStudy (anonymous):

It's ok. I tend to forget minor details when it's my own work. :) There's no pressure on me when I'm helping someone else, so I'm more detail-oriented.

OpenStudy (anonymous):

hahaha it happens. can i tell u the steps of solving this equation both ways after u explained it? just to make sure i got it 110%

OpenStudy (anonymous):

OK. (a+b)^2=25 (a+b) = -5 (take the square root of both sides) Since we know that (a+b)=-5, we can put it in the second equation. (-5)^3=?? =-125

OpenStudy (anonymous):

i meant could i explain it the way u told me, but that's great too. thank you

OpenStudy (anonymous):

was that the second way of solving this equation?

OpenStudy (anonymous):

That is how I explained it except I used words instead of numbers.

OpenStudy (anonymous):

i meant that i wanted to tell u how to solve it and have u check my work lol okok thank you! wait, i still don't know, is there a 2nd way of solving this? or just that one way that u showed me?

OpenStudy (anonymous):

I suppose you could use the guess-and-check method of solving it (where you just keep guessing random numbers until you get the answer), but that would take forever. And I solved it for you up there. :) The answer is -125.

OpenStudy (anonymous):

hahaha i guess that is a possible option, check and guess. we were just preparing for the psat and this was one of the qs i got wrong and the teacher told us we needed to solve it 2 ways.

OpenStudy (anonymous):

I can't think of an easy second way. Maybe someone else will come along and give you one. :) Sorry!

OpenStudy (anonymous):

oh it's completely alright, u've been a great help. I literally would not have gotten anywhere without ur help. thank you

OpenStudy (anonymous):

My pleasure. :)

OpenStudy (anonymous):

if i hv more q's can u please help again, if u can?

OpenStudy (anonymous):

Sure. :)

OpenStudy (anonymous):

ure the best! (:

OpenStudy (anonymous):

hey Becki?

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

\[\sqrt{x}-4 > 1\] how many positive integers (x) satisfy the inequality?

OpenStudy (anonymous):

i dont't understand wat steps i need to take to solve this

OpenStudy (anonymous):

wud i hv to isolate the variable first?

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

wait i wrote it wrong, its' radical 4-x then out of radical its greater than 1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so then i wud hv radical 3 is greater tan x?

OpenStudy (anonymous):

So is it\[\sqrt{4-x}>1\]

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

So square both sides to get rid of the radical.

OpenStudy (anonymous):

wud i bring over the x on the right side or keep it under radical?

OpenStudy (anonymous):

ohhhhh ok squaring, so then i wud hv x^2 - 8X +16 is greater than 1

OpenStudy (anonymous):

then subtract 16 from the other side so it wud be -15?

OpenStudy (anonymous):

No. When you square something that's under a radical, you just have what's under the radical. Think like this: \[\sqrt{2}\times \sqrt{2}=2\]

OpenStudy (anonymous):

so then it wud be \[\sqrt{4-x} \times \sqrt{4-x}\]

OpenStudy (anonymous):

yes....which is?

OpenStudy (anonymous):

my calculator says "ERROR" lol

OpenStudy (anonymous):

Yeah, you can't do it on a calculator. Remember, when you square the square root of something, you are left with the something.

OpenStudy (anonymous):

so it will just b 4-x?

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

Yes. :) so 4-1>1

OpenStudy (anonymous):

i get it, like from ur example

OpenStudy (anonymous):

omgosh ure a genius

OpenStudy (anonymous):

okok one sec let me solve

OpenStudy (anonymous):

sooooo, \[x <3\]

OpenStudy (anonymous):

and that means that there are only #'s left, which are 1 and 2 and so the answer is two positive integers

OpenStudy (anonymous):

You got it. :)

OpenStudy (anonymous):

ahh ure amazing, it's so easy with ur help!

OpenStudy (anonymous):

is this the only method of solving this equation?

OpenStudy (anonymous):

Guess and check is the only other way I can think of . :)

OpenStudy (anonymous):

probably the only other method for most of them.

OpenStudy (anonymous):

THANK YOU SO MUCH again lol :)

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

can u help with another? lol sorry!

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

In the sequence above, the first term is \[- \frac{ 3 }{ 16 }\] and each term after the first is -4 time the preceding term. wat is the result of dividing the 50th term by the 48th term?

OpenStudy (anonymous):

oh here's the sequence

OpenStudy (anonymous):

\[- \frac{ 3 }{ 16 }, \frac{ 3 }{ 4 }, -3, 12, -48,...\]

OpenStudy (anonymous):

Ooh. Now you're making me think! :) I like it.

OpenStudy (anonymous):

hahaha im sorry

OpenStudy (anonymous):

i just have absolutely no clue as to where to start with this equation

OpenStudy (anonymous):

No, it's fine. Give me a sec. I love puzzles. :)

OpenStudy (anonymous):

hehehe alright, take ur time! (:

OpenStudy (anonymous):

I know the answer off the top of my head, now I just need to figure out how to explain to you how to do it.

OpenStudy (anonymous):

that was so quick! lol ik the answer is 16

OpenStudy (anonymous):

Yup. And it's because you're multiplying by -4 each time, so -4x-4=16. The ratio between 50 and 48, because you're multiplying each by -4, is 16. Let's see....OH! I think i have it....just a sec.

OpenStudy (anonymous):

\[\frac{ n(4x4) }{ n }\] So, if you have a number n (that would be your 48th number in this case), and you're dividing it into that number nx-4x-4, wouldn't you get 16? It wouldn't really matter what numbers in the sequence you use as long as the numbers are 2 apart in the sequence. Does that make sense?

OpenStudy (anonymous):

it makes sense but let me just read it a few times to process lol one sec

OpenStudy (anonymous):

ok so i understand how they got the sequence (by x by -4), but i dont understand how we know what the 48the and 50th terms are..or is the 48th term the -48 in the sequence? so then do u multiply to get 192 and then multiply 192 by -4 = -768....wait no, no scratch that

OpenStudy (anonymous):

We don't need to know the 48th or 50th terms. It wouldn't matter. All that matters is we know that the top number is the bottom number times (-4) twice.

OpenStudy (anonymous):

Try it with any 2 other terms that are 2 places apart in the sequence. Try it with -3 and -48...

OpenStudy (anonymous):

And the question doesn't ask you to identify the 48th and 50th terms. It just asks you for the ratio.

OpenStudy (anonymous):

OHHHHH i see. when i divide -48 by -3 i get 16

OpenStudy (anonymous):

ohhh alright, cool

OpenStudy (anonymous):

other method of solving wud b guess and check again, right? haha

OpenStudy (anonymous):

Right?? I love numbers. They always do what they're supposed to do. :) Yeah, no other methods, once again. :) I have no idea what your teacher could be talking about.

OpenStudy (anonymous):

me either........lol

OpenStudy (anonymous):

thank you thank you thank you

OpenStudy (anonymous):

Again, you're welcome. I'm a total nerd. I got on here once for help in calculus and now I'm addicted to helping other people. I love this stuff.

OpenStudy (anonymous):

Come find me again sometime if you ever need help again. You're fun because you don't want me to feed you the answers like some that are here. :)

OpenStudy (anonymous):

heheheh yes, I'd really like to learn how to do them, i just need some guidance and u were really super helpful. thank so much! and i can c how it can become addicting, it's fun!